OpenStudy (anonymous):

Devon won a tiny bouncy ball at the school carnival. The diameter of the ball is 1.25 inches. To the nearest hundredth of a square inch, what is the surface area of the ball? Use 3.14 for pi.

OpenStudy (anonymous):

SA = 4πr^2 SA = 4π)(1.25/2)^2 SA = 4(π)(1.5625/4) SA = 1.5625π SA ≈ 4.91

OpenStudy (anonymous):

Your awesome! thanks a lot again! Wasn't sure if my procedure was correct!